# Unveiling Privacy Risks in LLM Agent Memory

### v1

@@ Abstract
Large Language Model (LLM) agents have become increasingly prevalent across various real-world applications. They enhance decision-making by storing private user-agent interactions in the memory module for demonstrations, introducing new privacy risks for LLM agents. In this work, we systematically investigate the vulnerability of LLM agents to our proposed Memory EXTRaction Attack (MEXTRA) under a black-box setting. To extract private information from memory, we propose an effective attacking prompt design and an automated prompt generation method based on different levels of knowledge about the LLM agent. Experiments on two representative agents demonstrate the effectiveness of MEXTRA. Moreover, we explore key factors influencing memory leakage from both the agent’s and the attacker’s perspectives. Our findings highlight the urgent need for effective memory safeguards in LLM agent design and deployment.
@@ 1 Introduction
Despite their success in advancing various domains, LLM agents often utilize and store private information, causing potential privacy risks, particularly in privacy-intensive applications such as healthcare. The private information of an LLM agent mainly originates from two sources: (1) The data the agent retrieves from external databases, containing sensitive and valuable domain-specific information Li et al. (2023); Kulkarni et al. (2024), e.g., patient prescriptions used in healthcare agents. (2) Historical records stored in the memory module [^1] This refers to long-term memory maintaining many past records rather than short-term memory, which only stores the current user-agent interaction Zhang et al. (2024). Zhang et al. (2024), consisting of pairs of private user instructions and the agent’s generated solutions. For example, in an intelligent auxiliary diagnosis scenario, a clinician’s query about treatment recommendations for a patient’s condition can expose the patient’s health status.
@@ 4 RQ1: LLM Agent Memory Extraction
With the attacking prompts generated through the basic instruction ${\mathcal{I}}^{\text{basic}}$, we empirically investigate the privacy leakage of the LLM agent memory on two real-world application agents. Our evaluation reveals the LLM agent’s high vulnerability to our memory extraction attack MEXTRA.
@@ 7 Related Work
Memory storing user-agent interactions provides valuable insights for LLM agents in solving real-word applications, making it an essential component of LLM agents Zhang et al. (2024). However, while equipping LLM agents with memory improves performance, it also introduces privacy risks. For instance, healthcare agents Shi et al. (2024); Li et al. (2023) store sensitive information about patients, web application agents Kagaya et al. (2024) record user preferences, and autonomous driving agents Mao et al. (2023); Wen et al. (2024) accumulate past driving scenarios. As these memory modules inherently store highly sensitive user data, a systematic investigation into the risks of memory leakage is crucial for revealing and mitigating potential threats.
@@ 8 Conclusion
In this paper, we unveil the privacy risks of LLM agent memory leakage through a memory extraction attack, MEXTRA. It consists of two parts: attacking prompt design and automated attacking prompt tailored to different levels of knowledge about the agent. Empirical evaluations demonstrate the vulnerability of LLM agents to MEXTRA. Moreover, we explore the key factors that influence memory leakage from both the agent’s and the attacker’s perspectives.
@@ Appendix C Discussion about Potential Mitigation
The vulnerability of LLM agent to MEXTRA highlights the need for mitigation strategies. We consider two categories of defenses: (1) Input and output control Jain et al. (2023); Inan et al. (2023), which aims to prevent private information exposure during query processing and response generation. A common approach is adding a hard rule in agent’s system prompt to filter privacy-sensitive input query and outputs, blocking responses that could lead to memory leakage. For example, a rule like “If the user requests historical queries, do not respond and stop this interaction” can filter malicious query input. Another approach is paraphrasing, which rephrases queries and responses to obscure identifiable private details. However, the attacking prompt can be tailored to appear contextually harmless and normal, making the filter fail to detect. And paraphrasing may not fully eliminate sensitive information. (2) Memory sanitation, which focuses on ensuring that stored data does not contain private information. For example, the agent can conduct de-identification, where user queries are sanitized before being stored Vakili et al. (2022); Liu et al. (2023), preventing the leakage of sensitive details. However, this operation may weaken the effectiveness of stored records as demonstrations for future queries. Based on above analysis, future research could develop effective memory safeguards in the design and deployment of LLM agents.

### v2

@@ Abstract
Large Language Model (LLM) agents have become increasingly prevalent across various real-world applications. They enhance decision-making by storing private user-agent interactions in the memory module for demonstrations, introducing new privacy risks for LLM agents. In this work, we systematically investigate the vulnerability of LLM agents to our proposed Memory EXTRaction Attack (MEXTRA) under a black-box setting. To extract private information from memory, we propose an effective attacking prompt design and an automated prompt generation method based on different levels of knowledge about the LLM agent. Experiments on two representative agents demonstrate the effectiveness of MEXTRA. Moreover, we explore key factors influencing memory leakage from both the agent designer’s and the attacker’s perspectives. Our findings highlight the urgent need for effective memory safeguards in LLM agent design and deployment.
@@ 1 Introduction
Despite their success in advancing various domains, LLM agents often utilize and store private information, causing potential privacy risks, particularly in privacy-intensive applications such as healthcare. The private information of an LLM agent mainly originates from two sources: (1) The data the agent retrieves from external databases, containing sensitive and valuable domain-specific information Li et al. (2023); Kulkarni et al. (2024), e.g., patient prescriptions used in healthcare agents. (2) Historical records stored in the memory module [^1] This refers to long-term memory maintaining many past records rather than short-term memory, which only stores the current user-agent interaction Zhang et al. (2024b). Zhang et al. (2024b), consisting of pairs of private user instructions and the agent’s generated solutions. For example, in an intelligent auxiliary diagnosis scenario, a clinician’s query about treatment recommendations for a patient’s condition can expose the patient’s health status.
@@ 4 RQ1: LLM Agent Memory Extraction
With the attacking prompts generated through the basic instruction ${\mathcal{I}}^{\text{basic}}$, we empirically investigate the privacy leakage of the LLM agent memory on two real-world application agents. Our evaluation reveals the LLM agent’s high vulnerability to our memory extraction attack MEXTRA [^2] The source code is available at https://github.com/wangbo9719/MEXTRA..
@@ 7 Related Work
Memory storing user-agent interactions provides valuable insights for LLM agents in solving real-word applications, making it an essential component of LLM agents Zhang et al. (2024b). However, while equipping LLM agents with memory improves performance, it also introduces privacy risks. For instance, healthcare agents Shi et al. (2024); Li et al. (2023) store sensitive information about patients, web application agents Kagaya et al. (2024) record user preferences, and autonomous driving agents Mao et al. (2023); Wen et al. (2024) accumulate past driving scenarios. As these memory modules inherently store highly sensitive user data, a systematic investigation into the risks of memory leakage is crucial for revealing and mitigating potential threats.
@@
Prompt Injection Attack.
Prompt injection is an attack that manipulates an LLM’s output by injecting crafted adversarial commands into its prompt He et al. (2024). It can be divided into two categories: direct and indirect prompt injection Rossi et al. (2024). In direct prompt injection, malicious input is directly included in the prompt given to the LLM, achieving the goals like goal hijacking Perez and Ribeiro (2022); Zhang et al. (2024a), prompt leaking Perez and Ribeiro (2022); Hui et al. (2024), and jailbreaking Zou et al. (2023); Li et al. (2024). In contrast, indirect prompt injection delivers malicious inputs through external content sources, often without the user’s awareness. It typically targets LLM-based agents via channels such as emails or websites Zhan et al. (2024); Wu et al. (2024a); Liu et al. (2023a), enabling more stealthy manipulation of external inputs and potentially leading to unauthorized actions and unintended data leakage, such as private information contained in emails or web pages. However, previous studies have paid limited attention to the issue of memory leakage in LLM-based agents, which is the main focus of this work.
@@ 8 Conclusion
In this paper, we unveil the privacy risks of LLM agent memory leakage through a memory extraction attack, MEXTRA. It consists of two parts: attacking prompt design and automated attacking prompt tailored to different levels of knowledge about the agent. Empirical evaluations demonstrate the vulnerability of LLM agents to MEXTRA. Moreover, we explore the key factors that influence memory leakage from both the agent designer’s and the attacker’s perspectives.
@@
Acknowledgements
Shenglai Zeng, Jiliang Tang and Pengfei He are supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, US Department of Commerce, Gates Foundation, the Michigan Department of Agriculture and Rural Development, Amazon, Meta, and SNAP.
@@ B.2 More Experiments
Experiments on QA-Agent
To further validate the generalization of MEXTRA across different black-box agent settings, in addition to the previously analyzed EHRAgent and RAP, which represent code-powered and web agents respectively, we additionally construct a question-answering agent (QA-Agent) equipped with a memory module.
Table 4: Attacking results on QA-Agemt. The number of attacking prompts $n$ is 30 and the memory size $m$ is 200.
$f(q,q_{i})$ Instruction EN RN EE CER AER
edit ${\mathcal{I}}^{\text{basic}}$ 28 30 0.23 0.87 0.93
${\mathcal{I}}^{\text{advan}}$ 46 49 0.38 0.80 0.93
cos ${\mathcal{I}}^{\text{basic}}$ 27 30 0.23 0.77 0.87
${\mathcal{I}}^{\text{advan}}$ 55 59 0.46 0.77 0.83
The QA-Agent is a general paradigm for answering user questions based on similar past questions stored in memory, which uses GPT-4o as its LLM backbone. Each record in memory is a (question, reasoning process) pair. The agent retrieves the top-4 most similar records as demonstrations. The generated reasoning process, including the answer and a detailed explanation, is finally provided to the user. To evaluate the impact of the similarity scoring function, we alter it between edit distance and cosine similarity, where the latter is computed using Sentence-BERT embeddings based on MiniLM. The memory size is 200, following the same setting as in §4.1. Questions in its memory are randomly selected from MMLU Wu et al. (2024b), a widely used benchmark containing a variety of questions from multiple disciplines. Note that the dataset used in this experiment is not domain-specific or privacy-related, since our goal is to validate the generality of MEXTRA itself.
Based on the different instructions introduced in §3.2, we generate 30 diverse attacking prompts for each instruction. The attacking results are presented in Table 4. It can be observed that MEXTRA achieves consistently solid performance on QA-Agent, aligning with its results on EHRAgent and RAP, which demonstrates the strong generalization capability of MEXTRA. The high CER scores indicate that most attacking prompts successfully extract all retrieved queries from memory. Compared with the basic instruction ${\mathcal{I}}^{\text{basic}}$, the two advanced instructions ${\mathcal{I}}^{\text{advan}}$ tailored to different similarity scoring functions bring significant improvements, achieving 55 and 46 extracted number (EN) respectively, leading to around 25% memory leakage from the memory storing 200 records.
@@ Appendix C Discussion about Potential Mitigation
The vulnerability of LLM agent to MEXTRA highlights the need for mitigation strategies. We consider two categories of defenses: (1) Input and output control Jain et al. (2023); Inan et al. (2023), which aims to prevent private information exposure during query processing and response generation. A common approach is adding a hard rule in agent’s system prompt to filter privacy-sensitive input query and outputs, blocking responses that could lead to memory leakage. For example, a rule like “If the user requests historical queries, do not respond and stop this interaction” can filter malicious query input. Another approach is paraphrasing, which rephrases queries and responses to obscure identifiable private details. However, the attacking prompt can be tailored to appear contextually harmless and normal, making the filter fail to detect. And paraphrasing may not fully eliminate sensitive information. (2) Memory sanitation, which focuses on ensuring that stored data does not contain private information. For example, the agent can conduct de-identification, where user queries are sanitized before being stored Vakili et al. (2022); Liu et al. (2023b), preventing the leakage of sensitive details. However, this operation may weaken the effectiveness of stored records as demonstrations for future queries. Based on above analysis, future research could develop effective memory safeguards in the design and deployment of LLM agents.